OpenStudy (anonymous):

A carpenter is building a rectangle room with a fixed perimeter of 172ft. What dimensions would yield the maximum area? What is the maximum area? The length that would yield the maximum area is ? ft

OpenStudy (anonymous):

I need the area please?

OpenStudy (anonymous):

well, say that the length of the area is x. then the width is (172 - 2x)/2, which is 81 - x so the area is x(81 - x) = 81x - x^2 to maximize the area, you find the turning point of A = -x^2 + 81x A = -(x^2 - 81x) A = -((x - 81/2)^2 - (81)^2/4) A = -(x - 81/2)^2 + (81)^2/4 so the maximum area is 81*81/4 = 6561/4 square feet

OpenStudy (anonymous):

and the length that would yield it is 81/2 feet

OpenStudy (anonymous):

1849 is what the answer was

OpenStudy (anonymous):

yes, that's because i'm an idiot who thought 172/2 was 86. my lord.

OpenStudy (anonymous):

er, was 81***

OpenStudy (anonymous):

but it's 86... so, call the length x, then the width is 86 - x thus the area is x(86 - x) = -x^2 + 86x A = -(x^2 - 86x) = -((x - 43)^2 - (43)^2) so A = -(x - 43)^2 + (43)^2 so maximum area is 43*43 = 1849 and length is 43

OpenStudy (anonymous):

width is 43 as well... it makes sense. the maximum area of a rectangle occurs when it is a square